OpenStudy (godlovesme):

HELPPPPPP ASAP PLEASE!! D: Using the following equation, find the center and radius of the circle. x2 + 2x + y2 + 4y = 20

OpenStudy (acxbox22):

sorry i gtg ^ that links explains what you have to do

OpenStudy (godlovesme):

awh man :c

OpenStudy (godlovesme):

its ok i'll try to figure it out thanks tho :)

OpenStudy (godlovesme):

@pooja195 help me please :((

OpenStudy (asnaseer):

you need to transform this equation:\[x^2 + 2x + y^2 + 4y = 20\]into the standard form:\[(x-h)^2+(y-k)^2=r^2\]which gives a circle of radius \(r\) and center at \((h,k)\)

OpenStudy (asnaseer):

start by looking at just the terms involving \(x\), i.e.:\[x^2+2x\]can you transform this into a form like \((x+?)^2\)?

OpenStudy (godlovesme):

i'll try, sorry im not good at this

OpenStudy (acxbox22):

im back

OpenStudy (asnaseer):

we know that:\[(x+a)^2=x^2+2ax+a^2\]

OpenStudy (asnaseer):

and we want to get:\[x^2+2x\]

OpenStudy (acxbox22):

well like the website says, you need to transform it into the center radius form by completing the square

OpenStudy (asnaseer):

can you see what value of \(a\) we need to pick?

OpenStudy (acxbox22):

nvr mind i will let @asnaseer explain :P

OpenStudy (godlovesme):

what do u mean @asnaseer

OpenStudy (godlovesme):

thanks though :)) @acxbox22

OpenStudy (asnaseer):

we know that:\[(x+a)^2=x^2+2ax+a^2\]can you see what value of \(a\) we need to pick so that we get a \(2x\) in the expansion on the right-hand-side?

OpenStudy (godlovesme):

is it 1?

OpenStudy (asnaseer):

correct! :)

OpenStudy (godlovesme):

phew than God i felt so stupid lol

OpenStudy (asnaseer):

so we know that:\[(x+1)^2=x^2+2x+1\]agreed so far?

OpenStudy (godlovesme):

yes :)

OpenStudy (asnaseer):

but we need just \(x^2+2x\), so we need to subtract 1 from both sides to get:\[(x+1)^2-1=x^2+2x\]agreed?

OpenStudy (asnaseer):

i.e. we know:\[(x+1)^2=x^2+2x+1\]\[\therefore (x+1)^2-1=x^2+2x+1-1=x^2+2x\]

OpenStudy (godlovesme):

ohhh so now we're left with \[x^2+2x\] ?

OpenStudy (asnaseer):

exactly

OpenStudy (asnaseer):

now remember your original equation was:\[x^2 + 2x + y^2 + 4y = 20\]we can now substitute \(x^2+2x\) with the new expression we found to get:\[(x+1)^2-1+y^2+4y=20\]understand?

OpenStudy (godlovesme):

yes :)

OpenStudy (asnaseer):

good, now lets first simplify this by adding 1 to both sides to get rid of the -1 from the left-hand-side

OpenStudy (asnaseer):

\[(x+1)^2-1+y^2+4y=20\]\[\therefore (x+1)^2-1+y^2+4y+1=20+1\]\[(x+1)^2+y^2+4y=21\]following so far?

OpenStudy (godlovesme):

yes :)

OpenStudy (asnaseer):

great! so now we have:\[(x+1)^2+y^2+4y=21\]lets now concentrate on the terms involving \(y\), we have:\[y^2+4y\]here again we know that:\[(y+b)^2=y^2+2by+b^2\]so can you think of what value we need for \(b\) so that the \(2by\) term becomes \(4y\)?

OpenStudy (godlovesme):

2 :3

OpenStudy (asnaseer):

perfect! :)

OpenStudy (asnaseer):

so we now have:\[(y+2)^2=y^2+4y+4\]but we are after just \(y^2+4y\), so we need to subtract 4 from both sides, i.e.:\[(y+2)^2=y^2+4y+4\]\[\therefore (y+2)^2-4=y^2+4y+4-4=y^2+4y\]agreed?

OpenStudy (godlovesme):

yes :D

OpenStudy (asnaseer):

ok, now recall we originally were given:\[x^2 + 2x + y^2 + 4y = 20\]which we transformed using \((x+1)^2-1=x^2+2x\) into:\[(x+1)^2+y^2+4y=21\]and now we also know that:\[(y+2)^2-4=y^2+4y\]so if we now replace \(y^2+4y\) with the new expression we just found for it, then we get:\[(x+1)^2+(y+2)^2-4=21\]agreed?

OpenStudy (godlovesme):

agreed c:

OpenStudy (asnaseer):

now again we can simplify this bu adding 4 to both sides as follows:\[(x+1)^2+(y+2)^2-4=21\]\[\therefore (x+1)^2+(y+2)^2-4+4=21+4\]\[\therefore (x+1)^2+(y+2)^2=25\]

OpenStudy (asnaseer):

agreed?

OpenStudy (godlovesme):

i get it now!! so the center will be (-1,-2) and the radius will be 5? :D

OpenStudy (asnaseer):

perfect! I was going to add one more step saying you have to be careful now as we are trying to get to this form:\[(x-h)^2+(y-k)^2=r^2\]so we must rewrite our equation as:\[(x-(-1))^2+(y-(-2))^2=5^2\]but I guess you beat me to it! :D

OpenStudy (godlovesme):

OMG THANK YOU SO MUCHHH!!!! you're a life saver, i really appreciate your help :))) <3

OpenStudy (asnaseer):

yw my friend. I really appreciate people who are genuinely interested in learning - keep up the good work! :)

OpenStudy (godlovesme):

ahh you're awesome!! I spend more than an hour trying to figure this out lol u solved it in less that half an hour XD

OpenStudy (asnaseer):

ok, bye for now - take care my friend :)

OpenStudy (godlovesme):

thx and you too :)